### v1

# Uncovering the Hidden Threat of Text Watermarking from Users with Cross-Lingual Knowledge

Uncovering the Hidden Threat of Text Watermarking from Users with Cross-Lingual Knowledge
Mansour Al Ghanim  Jiaqi Xue  Rochana Prih Hastuti Mengxin Zheng  Yan Solihin  Qian Lou
University of Central Florida
{mansour.alghanim,jiaqi.xue,rochana,mengxin.zheng,yan.solihin,qian.lou}@ucf.edu
In this study, we delve into the hidden threats posed to text watermarking by users with cross-lingual knowledge. While most research focuses on watermarking methods for English, there is a significant gap in evaluating these methods in cross-lingual contexts. This oversight neglects critical adversary scenarios involving cross-lingual users, creating uncertainty regarding the effectiveness of cross-lingual watermarking. We assess four watermarking techniques across four linguistically rich languages, examining watermark resilience and text quality across various parameters and attacks. Our focus is on a realistic scenario featuring adversaries with cross-lingual expertise, evaluating the adequacy of current watermarking methods against such challenges.
Uncovering the Hidden Threat of Text Watermarking from Users with Cross-Lingual Knowledge
Mansour Al Ghanim  Jiaqi Xue  Rochana Prih Hastuti Mengxin Zheng  Yan Solihin  Qian Lou University of Central Florida {mansour.alghanim,jiaqi.xue,rochana,mengxin.zheng,yan.solihin,qian.lou}@ucf.edu
1 Introduction
The advancement of Large Language Models (LLMs) has significantly transformed text generation across various domains, producing outputs that often closely mimic human writing. This advancement has raised concerns within the research community regarding potential misuses, including academic misconduct, the spread of disinformation, and the creation of synthetic training data (Xue et al., 2024b; Bender et al., 2021; Xue et al., 2024c; Zheng et al., 2024; Xue et al., 2024a; Cai et al., 2024). In response to these challenges, watermarking methods have been developed to differentiate between human-written and AI-generated texts (Aaronson, 2022; Kirchenbauer et al., 2023a; Kuditipudi et al., 2023; He et al., 2024; Dathathri et al., 2024; Chang et al., 2024).
Watermarking involves embedding a signal in AI-generated texts to identify the generating LLM using hypothesis testing. Specifically, watermarking offers theoretical guarantees regarding the detectability of the embedded signal by performing statistical inference on the generated text and testing against the null hypothesis. Since watermarking alters the generation of the original LLM, it is crucial to ensure that the impact on text quality is minimal while maintaining the watermark’s detectability. (1)(1)(1)Although steganography and watermarking share the practice of embedding signals, steganography primarily aims to conceal information through alterations of the text’s meaning, whereas watermarking serves to assert text source without changing the text’s semantics.
Much of the existing literature on watermarking Zhang et al. (2025) has primarily focused on the quality and detectability of watermarked texts, with a predominant emphasis on English texts. While the theoretical principles of these methods can be applied across languages, cross-lingual studies can reveal new adversarial scenarios in which watermark signal could be removed, and that have yet to be thoroughly investigated.
A recent work by He et al. (2024) addressed adversarial scenarios within a cross-lingual context, examining the consistency of watermarking through translation attacks, and developed defenses in this specific setting. Other existing studies on watermarking robustness have largely regarded back translation as the primary method for examining the interplay between languages, often overlooking other potential adversarial scenarios that may emerge following a translation to non-English languages  (Kuditipudi et al., 2023; Zhao et al., 2023; Pang et al., 2024). This oversight accounts for both unintentional alterations made by users for clarity or context adjustment and sophisticated adversaries intentionally refining the text to destroy watermark traces. In this paper, we evaluate four representative watermarking methods under two high-level themes. First, syntactical watermarking, which involves syntax changes to the generated text by manipulating the log-probabilities (logits) Lou et al. (2022) before the decode stage. Second, Semantic watermarking, which involves semantics manipulation of the generated text before the decode stage as well. The syntactical methods we consider are KGW  (Kirchenbauer et al., 2023a), EXP  (Aaronson, 2022), and Unigram  (Zhao et al., 2023). For the semantic methods, we choose XSIR  (He et al., 2024) as it aligns with our cross-lingual investigation in this paper. Specifically, this paper addresses the following research questions.
  • $\bullet$
    RQ1: How do watermarking methods perform across languages in terms of detectability, text quality, and diversity?
  • $\bullet$
    RQ2: How resilient are watermarking methods against existing cross-lingual attacks, particularly translation attacks?
  • $\bullet$
    RQ3: How do different adversarial approaches to cross-lingual watermark evasion perform? Specifically, how do attacks fare under our novel asymmetrical threat model, where the output language differs from the pivotal language?
Our investigation reveals that current watermarking methods face significant challenges in cross-lingual settings. While methods like XSIR shows promise for cross-lingual scenarios, all current approaches demonstrate vulnerabilities to different translation-based attacks. Furthermore, we find that traditional quality metrics may not adequately capture the diversity of cross-lingual watermarked text, necessitating new evaluation approaches. The main contributions of this paper are as follows:
  • $\bullet$
    We investigate the role of text watermarking across four different languages and four SOTA text watermarking methods, with a concentration on analyzing watermark detectability and quality.
  • $\bullet$
    We propose a new text diversity metric that utilizes Self-BLEU to effectively evaluate the quality of watermarked text generated in a cross-lingual context.
  • $\bullet$
    We evaluate the detectability of watermarks under practical attack pipeline consisting of translation, translation then paraphrase, and subsequent translation to the original language, and highlight the results on a robust cross-lingual method.
2 Background and Related Work
Syntactical-based Watermarking. Syntactical-based watermarking involves editing the generated text by manipulating the log-probabilities of a certain model. Multiple studies in the literature has investigated this type by introducing hypothesis testing. Pioneering in this area, the studies in KGW (Kirchenbauer et al., 2023a, b) and EXP (Aaronson, 2022) generate the watermark as a function of hashing the previous $k-1$-gram of a prompt to generate the next $k$ token. They introduce hypothesis testing to provide theoretical guarantees of watermark detectability. This is done by performing a statistical inference on the generated text, and calculating a score that is compared against a threshold with a very low False Positive Rate (FPR). Consequently, a barrage of similar watermark studies have been conducted to investigate the robustness and quality of KGW and EXP watermarks. In Unigram Zhao et al. (2023), the authors argue the robustness of the watermark to removal attacks could be mitigated by hashing a pre-determined key (uni-gram) that is used for all generations. Kuditipudi et al. (2023); Christ et al. (2024) introduced distortion-free watermarking in which the distribution of watermarked and unwatermarked text is the same. Similarly, Dathathri et al. (2024) introduces speculative sampling to generate watermarks at scale. Lu et al. (2024); Lee et al. (2023) investigated the role of token entropy to the watermark detectability and quality and proposes methods for generating and detecting the watermark.
Semantic Watermarking. Syntactical watermarking can still be compromised by paraphrasing and back-translation watermark removal attacks. In response to this, a number of studies have been conducted to evaluate watermarking methods in such scenarios. SIR (Liu et al., 2024) introduced semantic hashing of previous context instead of tokens to combat security and robustness attacks related to k-gram methods. Consequently, He et al. (2024) introduced XSIR, which is an extension of SIR with cross-lingual settings in which semantically cross-lingual words are clustered together. Unlike SIR, XSIR is able to mitigate translation attacks between different languages. Most recently, the work in  Chang et al. (2024) introduces the use of watermarking in a blackbox setting where users can semantically mark generated output of closed-source LLMs without accessing the log-probabilities for intellectual property reasons. Similarly, Hou et al. (2023, 2024) introduce sentence-level clustering to semantically mark the generated output.
Post-hoc Detection Approaches. In our evaluate, we evaluate proactive detection methods in which the watermark signal is embedded in the text during generation. However, multiple studies have investigated passive/discriminator methods for AI-generated text detection. Tian (2023); Mitchell et al. (2023); Gehrmann et al. (2019) use some statistical patterns in AI-generated text and trained discriminator models that differentiate human-written from AI-generated texts. Alshammari and Elleithy (2024) introduces a method to detect AI-generated text in Arabic language by leveraging language-specific diacritics. Abdelnabi and Fritz (2021) investigates adversarial text watermarking by modifying the inner workings of a transformer model to embed a hidden watermark.
3 Threat Model
In our threat model for text watermarking, we adopt a practical pipeline for potential watermark removal attacks, recognizing the various methods an adversary might employ. This pipeline includes stages such as simple translation, translation followed by paraphrasing, and a combination of translation, paraphrasing, and translating back again to the original language as seen in Figure 1. We consider a setting where LLMs are closed-source, in which case the model owners are the main victim of watermark removal attacks.
1 : Existing symmetric cross-lingual attacks in comparison to our asymmetric attacks. (a) In Symmetrical translation attacks, a pivot language is used to ❶ translate the prompt. Then the original language is obtained ❷ without further edits are considered. (b) In Asymmetrical translation, the user ❶ translates to a target language. The user can use directly or ❷ optionally edit the text, further attenuating the watermark signal. Stage ❸ is also optional, which restores some of the watermark signal.
Attacker Knowledge. We consider a multilingual adversary who may operate at varying levels of sophistication. The adversary might be naive, unaware of the presence of watermarks, or advanced, intentionally attempting to evade detection. They understand how to utilize technological tools that can alter text, potentially impacting the embedded watermarks. It’s crucial to consider that regular users might unintentionally remove watermarks by prompting an LLM directly in one language before translating and making textual edits as can be seen in Figure 1 (b) at stage ②. Ignoring this setting can mislead focus in the literature toward only intentional attacks utilizing pivotal languages as shown in Figure 1 (a) stage ②.
Attacker Capability. The adversary is equipped with the ability to access AI chat interfaces, through which they begin prompting and engaging with content. They can leverage translation APIs to translate texts into different languages, which may weaken or eliminate watermark signals. Moreover, they have the capability to use probably less sophisticated LLM to paraphrase the text, thereby further disrupting the watermark signal.
This comprehensive threat model highlights the challenges that the watermarked text goes through in a cross-lingual setting. Our threat model underscores the necessity of developing sophisticated watermarking methods that can endure these intricate and pipelined attacks.
4 Methodology
1.4 Quality Evaluation
Evaluating watermarked text quality is essential for assessing watermark methods. Most studies focus on English, so we introduce a pipeline for multi-lingual LLMs. Initially, we used perplexity (PPL), but it wasn’t effective across languages. PPL results can be shown in Appendix A, Figures 5 and  6. For better assessment, we use GPT-Judger from Singh et al. (2023) with an advanced OpenAI model. Due to k-gram repetitions in watermarked text, we also apply Self-BLEU to measure diversity and repetition. We combine these methods to create a metric that adjusts diversity more effectively than relying solely on Self-BLEU.
GPT-Judger. Using an external judge to examine the quality of text has proven to be useful  (Singh and Zou, 2023; Zheng et al., 2023), especially when text quality metrics like perplexity are unable to capture all aspects of text quality. GPT-Judger covers a range of criteria that reflect the essential parts of text quality. These categories include
Relevance to the prompt, Depth of detail, Clarity of writing, Coherence and logical flow, Originality and insight, Use of specific examples, and Accuracy of information.
Additionally, from GPT-judger, we could extract soft wins and highlight which quality criterion of the text is affected the most by the watermarking method.
Self-BLEU SB. SB Zhu et al. (2018) measures the diversity in a generated text by measuring how similar sentences are to each other in the same text. SB is calculated by matching a sentence (hypothesis) with all other sentences (references) in the same text for each sentence in the text. Then an average score is calculated and returned as the final score.
Adjusted Diversity AD. Because SB can be misleading in measuring the diversity of text –for example mixing tokens from different languages yields very low SB score, hence more diversity in text– we create a new diversity metric that reflects unwanted diversity or excessive repetition in text by leveraging the Judger’s coherency criterion scores as follow:
$\text{AD}=w\text{SB}+(1-w)(1-\text{NC})$ )1(
Where NC indicates Normalized Coherency score from GPT-Judger, $w$ is a weight between 0 and 1 that can be adjusted based on how much importance we want to give to each metric. The term $(1-\text{NC})$ inverts the coherency score from GPT-judger so that a low coherency score (indicating problematic text) contributes to a higher ”unrealistic diversity” in the text. In our experiment, we choose $w$ to be $0.3$ as SB doesn’t catch the semantic level of the text as GPT-Judger does, yet SB could give a subtle indication of repetition in the text.
2.4 Watermark Methods
KGW (Kirchenbauer et al., 2023a). The KGW method embeds a watermark signal in generated text by manipulating log-probabilities (logits) of next token. The vocabulary $v$ is divided into green and red lists based on a split ratio $\gamma$. A secret key $S_{k}$ and a hash of the previous $k-1$ token ids seed a pseudorandom generator to produce the next token $k$. The generating model is either limited to the green list (hard watermark) or biased toward it (soft watermark) by adding a small value $\delta$ to the logits. Soft watermarking manages low-entropy contexts with few green list options.
Unigram (Zhao et al., 2023). Unigram, like KGW, divides the vocabulary $v$ into green list $v_{g}$ and red list $v_{r}$. The key distinction is that in Unigram, these lists are consistent for all generated tokens over the course of the generation process, as it does not utilize hashing with previous token IDs. Instead, Unigram employs the sha256 hashing algorithm, using a secret key $S_{k}$ to partition the vocabulary into the green and red lists. Subsequently, it applies the soft-watermark technique to adjust the logit values. This approach is expected to reduce the effectiveness of watermark frequency counting removal attacks, since the hashing process does not incorporate previous token IDs. Both KGW and Unigram use the following equation to detect the watermark.
$z=(|s|_{G}-\gamma|s|)/\sqrt{|s|\gamma(1-\gamma)}$ )2(
where $|s|_{G}$ is the number of green tokens in the generated text, and $\gamma=\frac{|v_{G}|}{|v|}$.
EXP Aaronson (2022). EXP uses exponential minimum sampling, which is a variant of the Gumbel trick  Papandreou and Yuille (2011), to bias the distribution of the next token generation. Specifically, like KGW, EXP uses $k-1$ context-window to seed a pseudorandom generator (PRG) to generate the next token $k$. However, instead of using soft watermark or applying a $\delta$ value to the logits, EXP uses PRG to generate random numbers $r_{t,i}$, which is the same size as the vocabulary of the generating model. Then, at position $t$, token $i$ is sampled by maximizing the following quantity.
$\text{argmax}_{i}(r_{t,i}^{1/p_{t,i}})$
when $p_{t,i}$ is very small, token $i$ will only be chosen if $r_{t,i}$ is close to one, which is very unlikely to happen. In terms of randomness, this sampling method will return the same token every time the same $k-1$ context is used for the PRG. The watermark is then detected by the following equation.
$\sum_{t=1}^{n}log(\frac{1}{1-r_{t,i}})$ )3(
XSIR He et al. (2024). XSIR is designed to enhance cross-lingual watermarking by ensuring that semantically similar prefix texts receive similar logit biases. Instead of directly hashing token IDs, XSIR hashes a semantic chunk of the prefix text to generate the logit bias for the next token $k$. This approach ensures that different prefix texts with similar meanings produce comparable logit bias distributions:
$\text{Sim}(\Delta(x),\Delta(y))\approx\text{Sim}(E(x),E(y))$ )4(
where $E$ is a multilingual embedding model, and $\Delta$ is the function that determines the watermark logit bias distribution for all tokens in the vocabulary. By doing so, XSIR ensures that tokens predicted under similar semantic prefixes receive similar watermarking bias.
Besides, XSIR also enforces consistent biases across words within the same semantic cluster. Specifically, words that share the same meaning across different languages are assigned identical biases:
$C(i)=C(j)\Rightarrow\Delta_{C(i)}=\Delta_{C(j)}$ )5(
where $C(i)$ represents the cluster index of word $i$. As a result, if tokens $i$ and $j$ are semantically equivalent, they receive the same logit bias, preserving watermark consistency across translations.
2 : GPT-judge results. We compute the average of watermarked and unwatermarked scores for $500$ generations to find that the criterion the judge complains the most about was coherency decrease.
5 Experimental Setup
In all our experiments, we assess the strength of the watermark both before and after various attacks across four watermarking methods. Following previous work Kirchenbauer et al. (2023a); Zhao et al. (2023); Dathathri et al. (2024), we evaluate watermark detection using Receiver Operating Characteristic (ROC) curves, which illustrate the trade-off between the false positive rate (FPR) and false negative rate (FNR). FPR indicates that the generated text is falsely flagged as human written, which is more critical in this context. Although theoretical guarantees exist for the FPRs of the watermark methods we study, we adopt an empirical approach. Specifically, we automatically calculate a threshold based on the scores of unwatermarked text and then compare it against the scores of watermarked text. This method ensures that a fair and consistent threshold is applied, particularly in a multilingual setting. For further details, Appendix B reports additional performance metrics (TPR, TNR, FPR, FNR) derived from the thresholds used by the evaluated methods.
LLMs. We use several multi-billion parameter models to generate and evaluate watermarked texts. Due to the lack of availability of open-source LLMs that support all of the languages we used in our study, we used different models for different languages. For Chinese and Indonesian, we use the Sailor2 1B and 8B variants Sailor2 Team (2024) for text generation and perplexity evaluation(2)(2)(2)Perplexity results are provided in the appendices.. For Arabic, we generate text with the Jais family 6.7B model Inception (2024) and assess perplexity using Acegpt 7B Huang et al. (2023). English is generated on both models. We generate watermarked text in different hyper-parameter settings and different languages with an average of 10 GPU hours for KGW and Unigram, an average of 7 hours for XSIR, and 3 hours for EXP. We also use OpenAI GPT-4o-mini-2024-07-18 as the GPT-based judger which is slower but better than GPT-3.5-turbo. Each experiment took around 20 minutes for 500 generations to assess text quality with GPT-Judger. For translation attacks, we employ OPUS-MT models Tiedemann and Thottingal (2020); Tiedemann et al. (2022). To ensure consistency with watermark protocols (e.g., using the same generation vocabulary for detection), we generate English texts with both the Jais and Sailor2 models; translations (for Chinese, Indonesian, or Arabic) are derived from the original English versions generated by their respective LLMs.
Watermark Parameters. For KGW Kirchenbauer et al. (2023a) and Unigram Zhao et al. (2023) we use three green list ratios, $\gamma\in\{0.1,0.5,0.9\}$, and three bias values, $\delta\in\{2,5,10\}$. For KGW, we also use a lefthash with a context size of 1(3)(3)(3)Distinct from Unigram, where a key is hashed once and reused for subsequent generations.. When applicable, we follow previous studies by using a z-score threshold of $4$. For EXP Aaronson (2022), we set a context length of $4$ for the hashing and consider text watermarked if the $p$-value falls below $10^{-4}$. For XSIR He et al. (2024), approximately half the vocabulary is split into green and red lists, i.e., $\gamma=0.5$. Although the original XSIR uses a bias value of $\delta=1$, we vary the $\delta\in\{2,5,10\}$ to facilitate extensive evaluation; we also adopt their watermark threshold of $0.2$.
Unless stated otherwise, all experiments use $\gamma=0.5$ and $\delta=2.0$ for KGW, Unigram, and XSIR, as these values empirically provide the best balance between watermark strength and text quality. Additional results with varying parameters are presented in the appendices, which confirm with previous analyses.
Dataset. For all experiments, we use $500$ examples from the C4 dataset Raffel et al. (2020), which is available in all evaluated languages. For English, we follow previous work by using the RealNewsLike split; for other languages, we use the main training split, as RealNewsLike is available only for English(4)(4)(4)https://huggingface.co.
6 Results
We perform experiments to answer the following research questions:
RQ1: How do watermarking methods perform across languages in terms of detectability, text quality, and diversity?
RQ2: How resilient are watermarking methods against existing cross-lingual attacks, particularly translation attacks?
RQ3: How do different adversarial approaches to cross-lingual watermark evasion perform? Specifically, how do attacks fare under our novel asymmetrical threat model, where the output language differs from the pivotal language?
1.6 RQ1: Watermarking Performance under Cross-lingual Setting
Quality. We evaluate text quality through two complementary approaches: GPT-Judger and our proposed diversity metrics. This dual evaluation provides a comprehensive view of how watermarking affects text quality across languages.
Figure 2 shows the result of obtaining the decreases of each criterion score after watermarking, indicating the text coherency is the most affected aspect in all languages. Moreover, XSIR and EXP showcase the largest degrade of the quality. Our analysis indicates that GPT-Judger frequently flags issues related to clarity of writing, including repetition and the overlapping of different language characters or unknown symbols.
1 : GPT-Judger Final Verdict Analysis. A soft-win is recorded when the watermarked text is judged to be of equal (Tie) or superior quality (Hard-win) compared to the non-watermarked version, reflecting the method’s ability to preserve text quality. Higher soft-win rates indicate better quality retention after watermarking.
Method Language Hard-Win $\uparrow$ Tie $\uparrow$ Soft-Win $\uparrow$
KGW English 0.42 0.05 0.47
Arabic 0.41 0.15 0.56
Chinese 0.31 0.26 0.57
Indonesian 0.29 0.18 0.47
Unigram English 0.38 0.08 0.46
Arabic 0.37 0.12 0.49
Chinese 0.24 0.27 0.51
Indonesian 0.27 0.14 0.41
XSIR English 0.22 0.12 0.34
Arabic 0.20 0.06 0.26
Chinese 0.11 0.24 0.35
Indonesian 0.16 0.12 0.28
EXP English 0.10 0.04 0.14
Arabic 0.14 0.10 0.25
Chinese 0.07 0.24 0.31
Indonesian 0.15 0.12 0.26
(a) KGW
(b) Unigram
(c) XSIR
(d) EXP
3 : Watermark detection ROC curves with AUC before attacks. We fix $\gamma=0.5$ and $\delta=2.0$ for KGW, Unigram, and XSIR. The watermark threshold is calculated automatically by comparing unwatermarked and watermarked scores for $500$ generations.
To further analyze the GPT-Judger results, we calculate soft-win rates in Table 1. A soft-win occurs when the judge favors the watermarked text or issues a tie. Among all methods, KGW achieves the highest soft-win rate. Nonetheless, the GPT-Judger occasionally struggles to deliver definitive judgments, particularly for non-English texts, as indicated by the high tie rates compared to English. To objectively capture this impact, we introduce our new metric rooted in the Self-BLEU (SB) metric.
2 : SelfBleu results for all watermark methods. For KGW, Unigram and XSIR, $\gamma=0.5$ and $\delta=2.0$. We consider this setting is the best for these methods with different languages. More results with varying the $\gamma$ and $\delta$ values in are in Appendix
Method Language self-bleu ($\downarrow$ more diverse) AD ($\downarrow$ better)
watermarked unwatermarked watermarked unwatermarked
KGW English 0.16 0.16 0.38 0.34
Arabic 0.11 0.12 0.36 0.35
Chinese 0.04 0.04 0.44 0.40
Indonesian 0.10 0.10 0.43 0.38
Unigram English 0.23 0.17 0.40 0.35
Arabic 0.16 0.12 0.41 0.35
Chinese 0.02 0.04 0.47 0.40
Indonesian 0.12 0.10 0.46 0.37
XSIR English 0.20 0.17 0.49 0.32
Arabic 0.14 0.12 0.45 0.33
Chinese 0.03 0.04 0.55 0.38
Indonesian 0.12 0.11 0.53 0.37
EXP English 0.19 0.17 0.57 0.29
Arabic 0.20 0.12 0.55 0.31
Chinese 0.13 0.04 0.61 0.38
Indonesian 0.21 0.10 0.57 0.36
4 : Watermark detection ROC curves with AUC after attacks. We fix $\gamma=0.5$ and $\delta=2.0$ for KGW, Unigram, and XSIR. For all attacks, the watermark threshold is calculated automatically by comparing unwatermarked and attacked watermarked text score for $100$ generations. Left: we perform translation attacks on KGW, Unigram, and EXP. Right: Only the leftmost column represent asymmetrical translation attack setting. We apply more invasive attacks to XSIR to highlight its robustness
Diversity. Previous research has used the Self-BLEU (SB) metric to assess the diversity of watermarked text (Dathathri et al., 2024). Table 2 includes results from SB and the Adjusted Diversity (AD) metrics for various watermark methods. While SB suggests high diversity, especially for Chinese and Indonesian, it can be misleading; low SB scores (e.g., $0.04$ for Chinese) might indicate text degradation, as shown by higher AD scores ($\geq 0.44$). This pattern holds across other languages. When we employ larger $\delta$ values with $\gamma=0.5$, AD ratios are higher compared to SB, indicating unrealistic diversity in watermarked text. More detailed results with different hyper-parameters are in Table 3 in Appendix A(5)(5)(5)Language mixing can be due to combined-language prompts, model size, or vocabulary coverage, and watermarking exacerbate the problem.
Detectability. Our detection analysis examines both performance without attack and attack resilience across languages. Figures 3 and 4 show the ROC curves for all watermarking methods before and after attacks. Prior to any attacks, all methods demonstrate strong detectability. For KGW, Unigram, and XSIR, we identified $\delta=2.0$ and $\gamma=0.5$ as the optimal parameters yielding the best detection results. However, the role of employing different hyper-parameters for detection differs from that for text quality. When larger values of $\gamma$ are used with $delta=2.0$, detection scores are adversely affected, especially for Unigram method. This is counterintuitive to the green/red splitting methodology as in KGW and Unigram since using $delta=2.0$ with smaller $\gamma=0.1$ yield higher detection scores. We show detailed results in Figure 7 in which we illustrate the close relationship between $\gamma$ and $\delta$ for methods like KGW and Unigram in Appendix B.
2.6 RQ2: Watermarking Resiliencies to Translation Attacks.
Regarding attack resilience, all methods vulnerable to our proposed asymmetrical attacks. In Figure 4 (left), we perform asymmetrical translation attacks against KGW, Unigram, and EXP. The results showcase the watermark detectability under translation attacks varies by language Ghanim et al. (2024). For instance, KGW shows the lowest AUC of 0.55 for Arabic-English, while Unigram presents a worst-case AUC of 0.57 for English-Arabic and Arabic-English. In the EXP method as shown in Figure 4 (c), all attacks are notably invasive, likely due to its larger hashing window $4$ compared to KGW’s recommended window of $1$.
In Figure 4 (right), we evaluate all of our asymmetrical attacks to the more robust XSIR method. Although XSIR was developed for symmetrical attacks, it remains susceptible to asymmetrical attacks. This vulnerability underscores the severity of attack scenarios where the pivot language differs from the output language. XSIR’s semantic clusters vary across languages, weakening its watermark robustness when the pivot and output languages differ. In contrast, back-translate attacks preserve the watermark better, as translating back to the pivot language realigns the semantic clusters as shown in Figure 4 (f).
Additionally, KGW and EXP have more predictable results than XSIR and Unigram where different languages affect the watermark signal differently. For instance, detection outcomes for translation and translation-paraphrase attacks remain consistently tight across languages for KGW and EXP, whereas XSIR and Unigram exhibit greater variability, with some languages deviating significantly. This behavior may stem from technical challenges such as preserving text length during translation.
3.6 RQ3: Asymmetrical Translation Attacks Raise a New Threat
Despite XSIR being specifically designed for cross-lingual scenarios, our results shown in Figure 4 demonstrate that it remains vulnerable to post-generation translation attacks. Particularly, the vulnerability is pronounced for Chinese and Arabic languages, suggesting that XSIR’s effectiveness varies significantly by language. This contrasts with XSIR proposed adversarial scenario of translating prompts both before and after generation, indicating that post-generation attacks can be more effective at evading detection.
Our analysis reveals a critical asymmetry: translations TO English yield better detection rates than translations FROM English. This asymmetry appears to be rooted in XSIR’s semantic clustering characteristics, as we found that XSIR semantic clusters have more similar English words than other languages. This disparity is likely due to XSIR’s semantic clustering, which groups English words more consistently than words from other languages. Consequently, XSIR appears to perform more robustly when English is the target language rather than the source language.
These findings highlight the need for more robust cross-lingual watermarking approaches that can maintain effectiveness across different languages and resist sophisticated attack pipelines.
7 Conclusion
In this paper, we assess four distinct watermarking techniques across four different languages, considering realistic watermark removal attacks in a cross-lingual context. We examine the effectiveness of these watermarking methods by evaluating the quality of the watermarked text and the resilience of the watermark signal following such attacks. Our findings indicate that various watermarking techniques respond differently to language translations when the watermarked text is modified. We propose that there is a need for more robust cross-lingual watermarking solutions that integrate the predictability of KGW and EXP across languages with the more resilient approaches of Unigram and XSIR.
8 Limitations and Discussions
In this study, we evaluate four distinct watermarking methods across four languages, focusing on practical removal attacks in a cross-lingual context. We believe that more research should include languages with rich syntax and structure and suggest exploring datasets beyond C4, given its generalized nature.
The attacks presented are invasive, significantly weakening the watermark signal, but can be countered through various strategies. We believe that developing more resilient cross-lingual watermarking techniques should integrate KGW and EXP predictability with the robust Unigram method. XSIR could be instrumental here, as it manipulates text semantics over syntax. We first analyze why XSIR falls short against translation attacks and then explore potential solutions.
Our investigation into XSIR revealed that its shortcomings in handling translation attacks are largely technical. XSIR clusters semantically similar words across languages by constructing a graph from a comprehensive dictionary using text transformations. It identifies connected components (CCs), but this process can be problematic for languages with diverse character sets. For instance, CCs rely on exact matches, which may prevent clustering of semantically similar words.
To enhance XSIR’s robustness, we suggest modifying the detection process. Given that English handles XSIR cluster transformations effectively, translating non-English outputs to English before detection could ensure watermark signal identification. This semantic-focused approach can extend to other methods, such as PostMark  (Chang et al., 2024).
Additionally, XSIR’s robustness can be improved by refining the use of CCs. Changing node content to latent word representations, formulated as numerical vectors, would enable more meaningful exact matching, enhancing clustering accuracy across different languages.
9 Ethics Consideration
Our research reveals critical vulnerabilities in text watermarking when subjected to cross-lingual translation attacks—specifically, the risk that the original language of a text can be concealed, allowing adversaries to evade detection and potentially disseminate harmful content. We acknowledge that such findings may be exploited by malicious actors, thereby posing a serious risk to digital authenticity and safety. However, the primary goal of our work is to illuminate these weaknesses so that more robust watermarking strategies can be developed and integrated into language models. In the interim, we propose simple yet effective countermeasures that can be readily incorporated by AI service providers. Our study employs publicly available data and models and is intended solely for academic research and the improvement of digital security. We strongly advocate for responsible disclosure and the continuous refinement of safeguards to ensure that AI technologies are deployed safely and ethically.
  • Aaronson (2022) Scott Aaronson. 2022. My AI Safety Lecture for UT Effective Altruism.
  • Abdelnabi and Fritz (2021) Sahar Abdelnabi and Mario Fritz. 2021. Adversarial watermarking transformer: Towards tracing text provenance with data hiding. In 2021 IEEE Symposium on Security and Privacy (SP), pages 121–140. IEEE.
  • Alshammari and Elleithy (2024) Hamed Alshammari and Khaled Elleithy. 2024. Toward robust arabic ai-generated text detection: Tackling diacritics challenges. Information, 15(7):419.
  • Bender et al. (2021) Emily M Bender, Timnit Gebru, Angelina McMillan-Major, and Shmargaret Shmitchell. 2021. On the dangers of stochastic parrots: Can language models be too big? In Proceedings of the 2021 ACM conference on fairness, accountability, and transparency, pages 610–623.
  • Cai et al. (2024) Kunbei Cai, Zhenkai Zhang, Qian Lou, and Fan Yao. 2024. Wbp: Training-time backdoor attacks through hardware-based weight bit poisoning. In European Conference on Computer Vision, pages 179–197. Springer.
  • Chang et al. (2024) Yapei Chang, Kalpesh Krishna, Amir Houmansadr, John Wieting, and Mohit Iyyer. 2024. Postmark: A robust blackbox watermark for large language models. arXiv preprint arXiv:2406.14517.
  • Christ et al. (2024) Miranda Christ, Sam Gunn, and Or Zamir. 2024. Undetectable watermarks for language models. In The Thirty Seventh Annual Conference on Learning Theory, pages 1125–1139. PMLR.
  • Dathathri et al. (2024) Sumanth Dathathri, Abigail See, Sumedh Ghaisas, Po-Sen Huang, Rob McAdam, Johannes Welbl, Vandana Bachani, Alex Kaskasoli, Robert Stanforth, Tatiana Matejovicova, et al. 2024. Scalable watermarking for identifying large language model outputs. Nature, 634(8035):818–823.
  • Gehrmann et al. (2019) Sebastian Gehrmann, Hendrik Strobelt, and Alexander M Rush. 2019. Gltr: Statistical detection and visualization of generated text. arXiv preprint arXiv:1906.04043.
  • Ghanim et al. (2024) Mansour Ghanim, Saleh Almohaimeed, Mengxin Zheng, Yan Solihin, and Qian Lou. 2024. Jailbreaking llms with arabic transliteration and arabizi. In Proceedings of the 2024 Conference on Empirical Methods in Natural Language Processing, pages 18584–18600.
  • He et al. (2024) Zhiwei He, Binglin Zhou, Hongkun Hao, Aiwei Liu, Xing Wang, Zhaopeng Tu, Zhuosheng Zhang, and Rui Wang. 2024. Can watermarks survive translation? on the cross-lingual consistency of text watermark for large language models. arXiv preprint arXiv:2402.14007.
  • Hou et al. (2023) Abe Bohan Hou, Jingyu Zhang, Tianxing He, Yichen Wang, Yung-Sung Chuang, Hongwei Wang, Lingfeng Shen, Benjamin Van Durme, Daniel Khashabi, and Yulia Tsvetkov. 2023. Semstamp: A semantic watermark with paraphrastic robustness for text generation. arXiv preprint arXiv:2310.03991.
  • Hou et al. (2024) Abe Bohan Hou, Jingyu Zhang, Yichen Wang, Daniel Khashabi, and Tianxing He. 2024. k-semstamp: A clustering-based semantic watermark for detection of machine-generated text. arXiv preprint arXiv:2402.11399.
  • Huang et al. (2023) Huang Huang, Fei Yu, Jianqing Zhu, Xuening Sun, Hao Cheng, Dingjie Song, Zhihong Chen, Abdulmohsen Alharthi, Bang An, Ziche Liu, Zhiyi Zhang, Junying Chen, Jianquan Li, Benyou Wang, Lian Zhang, Ruoyu Sun, Xiang Wan, Haizhou Li, and Jinchao Xu. 2023. Acegpt, localizing large language models in arabic. Preprint, arXiv:2309.12053.
  • Inception (2024) Inception. 2024. Jais family model card.
  • Kirchenbauer et al. (2023a) John Kirchenbauer, Jonas Geiping, Yuxin Wen, Jonathan Katz, Ian Miers, and Tom Goldstein. 2023a. A watermark for large language models. In International Conference on Machine Learning, pages 17061–17084. PMLR.
  • Kirchenbauer et al. (2023b) John Kirchenbauer, Jonas Geiping, Yuxin Wen, Manli Shu, Khalid Saifullah, Kezhi Kong, Kasun Fernando, Aniruddha Saha, Micah Goldblum, and Tom Goldstein. 2023b. On the reliability of watermarks for large language models. arXiv preprint arXiv:2306.04634.
  • Kuditipudi et al. (2023) Rohith Kuditipudi, John Thickstun, Tatsunori Hashimoto, and Percy Liang. 2023. Robust distortion-free watermarks for language models. arXiv preprint arXiv:2307.15593.
  • Lee et al. (2023) Taehyun Lee, Seokhee Hong, Jaewoo Ahn, Ilgee Hong, Hwaran Lee, Sangdoo Yun, Jamin Shin, and Gunhee Kim. 2023. Who wrote this code? watermarking for code generation. arXiv preprint arXiv:2305.15060.
  • Liu et al. (2024) Aiwei Liu, Leyi Pan, Xuming Hu, Shiao Meng, and Lijie Wen. 2024. A semantic invariant robust watermark for large language models. In The Twelfth International Conference on Learning Representations.
  • Lou et al. (2022) Qian Lou, Yepeng Liu, and Bo Feng. 2022. Trojtext: Test-time invisible textual trojan insertion. In The Eleventh International Conference on Learning Representations.
  • Lu et al. (2024) Yijian Lu, Aiwei Liu, Dianzhi Yu, Jingjing Li, and Irwin King. 2024. An entropy-based text watermarking detection method. arXiv preprint arXiv:2403.13485.
  • Mitchell et al. (2023) Eric Mitchell, Yoonho Lee, Alexander Khazatsky, Christopher D. Manning, and Chelsea Finn. 2023. DetectGPT: Zero-Shot Machine-Generated Text Detection using Probability Curvature.
  • Pang et al. (2024) Qi Pang, Shengyuan Hu, Wenting Zheng, and Virginia Smith. 2024. No free lunch in llm watermarking: Trade-offs in watermarking design choices. In The Thirty-eighth Annual Conference on Neural Information Processing Systems.
  • Papandreou and Yuille (2011) George Papandreou and Alan L Yuille. 2011. Perturb-and-map random fields: Using discrete optimization to learn and sample from energy models. In 2011 international conference on computer vision, pages 193–200. IEEE.
  • Raffel et al. (2020) Colin Raffel, Noam Shazeer, Adam Roberts, Katherine Lee, Sharan Narang, Michael Matena, Yanqi Zhou, Wei Li, and Peter J. Liu. 2020. Exploring the limits of transfer learning with a unified text-to-text transformer. Journal of Machine Learning Research, 21(140):1–67.
  • Sailor2 Team (2024) Sailor2 Team. 2024. Sailor2: Sailing in south-east asia with inclusive multilingual llm.
  • Singh and Zou (2023) Karanpartap Singh and James Zou. 2023. New evaluation metrics capture quality degradation due to llm watermarking. Transactions on Machine Learning Research.
  • Tian (2023) Edward Tian. 2023. Gptzero update v1.
  • Tiedemann et al. (2022) Jörg Tiedemann, Mikko Aulamo, Daria Bakshandaeva, Michele Boggia, Stig-Arne Grönroos, Tommi Nieminen, Alessandro Raganato, Yves Scherrer, Raul Vazquez, and Sami Virpioja. 2022. Democratizing machine translation with opus-mt. arXiv preprint arXiv:2212.01936.
  • Tiedemann and Thottingal (2020) Jörg Tiedemann and Santhosh Thottingal. 2020. OPUS-MT â Building open translation services for the World. In Proceedings of the 22nd Annual Conference of the European Association for Machine Translation.
  • Xue et al. (2024a) Jiaqi Xue, Qian Lou, and Mengxin Zheng. 2024a. Badfair: Backdoored fairness attacks with group-conditioned triggers. In Findings of the Association for Computational Linguistics: EMNLP 2024, pages 8257–8270.
  • Xue et al. (2024b) Jiaqi Xue, Mengxin Zheng, Yebowen Hu, Fei Liu, Xun Chen, and Qian Lou. 2024b. Badrag: Identifying vulnerabilities in retrieval augmented generation of large language models. arXiv preprint arXiv:2406.00083.
  • Xue et al. (2024c) Jiaqi Xue, Mengxin Zheng, Ting Hua, Yilin Shen, Yepeng Liu, Ladislau Bölöni, and Qian Lou. 2024c. Trojllm: A black-box trojan prompt attack on large language models. Advances in Neural Information Processing Systems, 36.
  • Zhang et al. (2025) Hang Zhang, Qian Lou, and Yanshan Wang. 2025. Towards safe ai clinicians: A comprehensive study on large language model jailbreaking in healthcare. arXiv preprint arXiv:2501.18632.
  • Zhao et al. (2023) Xuandong Zhao, Prabhanjan Ananth, Lei Li, and Yu-Xiang Wang. 2023. Provable robust watermarking for ai-generated text. arXiv preprint arXiv:2306.17439.
  • Zheng et al. (2023) Lianmin Zheng, Wei-Lin Chiang, Ying Sheng, Siyuan Zhuang, Zhanghao Wu, Yonghao Zhuang, Zi Lin, Zhuohan Li, Dacheng Li, Eric Xing, et al. 2023. Judging llm-as-a-judge with mt-bench and chatbot arena. Advances in Neural Information Processing Systems, 36:46595–46623.
  • Zheng et al. (2024) Mengxin Zheng, Jiaqi Xue, Xun Chen, Yanshan Wang, Qian Lou, and Lei Jiang. 2024. Trojfsp: Trojan insertion in few-shot prompt tuning. In Proceedings of the 2024 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies (Volume 1: Long Papers), pages 1141–1151.
  • Zhu et al. (2018) Yaoming Zhu, Sidi Lu, Lei Zheng, Jiaxian Guo, Weinan Zhang, Jun Wang, and Yong Yu. 2018. Texygen: A benchmarking platform for text generation models. SIGIR.
A More Quality Results
1.1 Perplexity (PPL)
Perplexity is a metric that, although known for its limitations, still provides a useful statistical or probabilistic background for assessing text quality. It measures how well a probability distribution predicts a sample. In the context of language models, it provides insights into how surprised the model is by the actual sequence of words compared to its predicted probabilities. A lower perplexity score indicates that the model found the text more predictable and thus, in a broad sense, of higher quality. In our experiments, the PPL is calculated as follows:
$\text{Perplexity}(W)=\exp\left(\frac{1}{N}\sum_{i=1}^{N}-\log p(w_{i})\right)$
where $\exp$ represents the exponential function. $N$ is the total number of words in the text. $p(w_{i})$ is the probability assigned to the word $w_{i}$ by the model. The negative log likelihood, $-\log p(w_{i})$, is referred to the Cross-Entropy loss for word $w_{i}$.
PPL Analysis:
(a) KGW
(b) Unigram
5 : Detection scores as a function of PPL for KGW and Unigram. For KGW, we notice that as $\delta$ grows higher, the quality of text decreases for all languages. Larger $\gamma$ values with smaller $\delta$ values greately affected the watermark strength in which it is attenuated. Unigram presents interesting graphs. When $\gamma$ is large, we see simiar trend as in KGW. However, smaller $\gamma$ values behave differently for different languages. The PPL and Z-scores are calculated on 500 generations.
(a) XSIR
(b) EXP
6 : Detection scores as a function of PPL for XSIR and EXP. XSIR employs $\gamma=0.5$. The smaller the $\delta$ the better the qulaity of text. EXP $p$-values show insensitivity to PPL scores with a few examples falling above the threshold or indicating False Negatives.
Figures  5 and  6 show the strength of detection of the watermark as a function of PPL. The results in figure  5(a) confirm the findings in  Kirchenbauer et al. (2023a) about the slight effect of larger $\delta$ values to the quality of the watermarked text for all languages. We see similar trends for Unigram in figure  5(b) only when $\gamma\geq 0.5$. For smaller values of $\gamma$, the quality of text seems to slightly improve even for larger $\delta$ values. We believe that this effect is a result of low-entropy completions from the red list being repeated over and over given smaller potion of the vocabulary, which increases the z-scores and precludes PPL from catching this effect.
For the other watermark methods XSIR and EXP, we show their results in figure  6. For EXP, we experiment with the whole data points to clearly identify the trends. In figure  6(a), XSIR shows similar trends to KGW in terms of the effect of larger values of $\delta$ on text quality. For EXP as shown in figure  6(b), the quality of text has a consistent relationship with the lower $p$-values across all languages. In other words, the $p$-values remain virtually insensitive to variations in the quality of the text.
2.1 Complete Self-BLEU Results
Table  3, we show complete results of self-BLEU and Adjusted Diversity (AD) using different hyper-parameter settings. From the results, it’s clear that when the $\delta$ is large, the text diversity unrealistically increases for KGW, Unigram and XSIR. This is due to the tight relationship between $\gamma$ and $\delta$ in which lower $\gamma$ values with higher $\delta$ values causes an adverse effect on the text quality.
Method Language $\gamma$ $\delta$ self-bleu ($\downarrow$ more diverse) AD ($\downarrow$ more diverse)
watermarked unwatermarked watermarked unwatermarked
KGW English 0.5 2.0 0.16 0.16 0.38 0.34
5.0 0.15 0.17 0.46 0.32
10.0 0.15 0.17 0.50 0.31
Arabic 0.5 2.0 0.11 0.12 0.36 0.35
5.0 0.12 0.12 0.41 0.35
10.0 0.12 0.13 0.45 0.33
Chinese 0.5 2.0 0.04 0.04 0.44 0.40
5.0 0.04 0.04 0.46 0.41
10.0 0.04 0.04 0.48 0.40
Indonesian 0.5 2.0 0.10 0.10 0.43 0.38
5.0 0.08 0.10 0.53 0.35
10.0 0.08 0.11 0.60 0.33
Unigram English 0.5 2.0 0.23 0.17 0.40 0.35
5.0 0.26 0.17 0.54 0.32
10.0 0.27 0.17 0.59 0.31
Arabic 0.5 2.0 0.16 0.12 0.41 0.35
5.0 0.19 0.12 0.48 0.32
10.0 0.21 0.12 0.52 0.34
Chinese 0.5 2.0 0.02 0.04 0.47 0.40
5.0 0.02 0.04 0.52 0.39
10.0 0.03 0.03 0.54 0.39
Indonesian 0.5 2.0 0.12 0.10 0.46 0.37
5.0 0.13 0.10 0.57 0.34
10.0 0.13 0.11 0.63 0.33
XSIR English 0.5 2.0 0.20 0.17 0.49 0.32
5.0 0.22 0.17 0.59 0.30
10.0 0.22 0.17 0.63 0.29
Arabic 0.5 2.0 0.14 0.12 0.45 0.33
5.0 0.19 0.13 0.55 0.30
10.0 0.19 0.13 0.57 0.31
Chinese 0.5 2.0 0.03 0.04 0.55 0.38
5.0 0.04 0.04 0.61 0.36
10.0 0.03 0.04 0.62 0.36
Indonesian 0.5 2.0 0.12 0.10 0.53 0.37
5.0 0.10 0.10 0.66 0.32
10.0 0.10 0.10 0.68 0.31
EXP English - - 0.19 0.17 0.57 0.29
Arabic 0.20 0.12 0.55 0.31
Chinese 0.13 0.04 0.61 0.38
Indonesian 0.21 0.10 0.57 0.36
3 : Complete SelfBleu results with the Adjusted Diversity (AD) metric for all watermark methods. For KGW, Unigram, and XSIR, we fix $\gamma=0.5$ and vary $\delta=[2.0,5.0,10.0]$
3.1 Complete Soft-Win Results
In Table  4, we show more results across different hyper-parameter settings for KGW, Unigram and XSIR. The soft-win rates drastically decreases with lower $\gamma$ values and higher $\delta$ values.
Method Language $\gamma$ $\delta$ Soft Win Rate Language Avg. Method Avg.
KGW English 0.5 2.0 0.47 0.314 0.355
5.0 0.264
10.0 0.21
Arabic 0.5 2.0 0.56 0.417
5.0 0.418
10.0 0.314
Chinese 0.5 2.0 0.566 0.502
5.0 0.514
10.0 0.426
Indonesian 0.5 2.0 0.468 0.267
5.0 0.224
10.0 0.108
Unigram English 0.5 2.0 0.462 0.287 0.314
5.0 0.232
10.0 0.166
Arabic 0.5 2.0 0.494 0.352
5.0 0.292
10.0 0.27
Chinese 0.5 2.0 0.51 0.436
5.0 0.422
10.0 0.376
Indonesian 0.5 2.0 0.412 0.234
5.0 0.178
10.0 0.112
XSIR English 0.5 2.0 0.342 0.256 0.218
5.0 0.2
10.0 0.226
Arabic 0.5 2.0 0.26 0.183
5.0 0.136
10.0 0.1
Chinese 0.5 2.0 0.35 0.293
5.0 0.262
10.0 0.268
Indonesian 0.5 2.0 0.276 0.15
5.0 0.108
10.0 0.066
EXP English - - 0.142 - 0.24
Arabic 0.246
Chinese 0.31
Indonesian 0.262
4 : Soft Win Rates for Different Methods by Language and Hyper-parameters. Columns are added for averages across languages and methods.
B Detailed Detection Results
(a) KGW
(b) Unigram
7 : Here we fix $\delta$ at $2.0$ and vary $\gamma$ for KGW and Unigram with $\gamma=(0.1,0.5,0.9)$ and with a fixed $\delta=2.0$. The larger the $\gamma$ ratio, the worse the watermark detection. We also use smaller FPRs in the range $[0.1,0.15]$
(a) XSIR
(b) EXP
8 : Watermark detection for XSIR and EXP with lower values of FPRs, which are in the range $[0.1,0.15]$. For XSIR, lower values of $\delta$ results in lower TPRs at very low FPRs.
1.2 Watermark Detection with Multiple $\gamma$ Ratios for KGW and Unigram
KGW  Kirchenbauer et al. (2023a) as shown in  3(a) performs the best among all other watermark methods, while Unigram as shown in figure  3(b) performs the worst. XSIR and EXP as shown in figures  3(c) and  3(d) respectively show higher TPRs compared to KGW and Unigram. In all watermark methods, English language shows stable AUC curve in all FPRs in all watermark methods. For KGW and Unigram, increasing the value of $\gamma$ and decreasing the value of $\delta$ has adverse effect on the TPR characteristic, especially for languages other than English (see subsequent sections for roc curves with zoomed-in rates).
Kirchenbauer et al. (2023a) has investigated the tight relationship of $\gamma$ and $\delta$ in watermark strength by creating a lower bound on $\gamma$ with the spike entropy in the picture. However, we believe more investigation is needed in cross-lingual manner. For example, beside the effect of the sampling method (greedy or multinomial), what could be done to incorporate the role of the tokenization of different languages in the $\gamma$ lower bound analysis? By doing a simple analysis of the results with $\gamma=0.9$ and $\delta=0.2$, we found that only an average of $2$ greenlist tokens are missing to raise the z-score from, say $3.52$ to $4$, with a z-score of threshold $4$. As shown in Figures 7(a) and 7(b), the z-score threshold performs adequately for both methods in English but struggles with other languages. At FPR 0%, KGW experiences the most significant TPR drop in Arabic and Indonesian, while Unigram reach worst TPR primarily in Chinese and Indonesian. This suggests that the current z-score setting may not be optimal for all languages, particularly for this $\gamma$ configuration. Further investigation into language-specific z-score adjustments is left for future work.
While this is more evident in KGW and Unigram, XSIR as shown in Figure 8(a) has relatively minor effects for smaller values of $\delta$. This is because XSIR logically uses $50\%$ of the total vocabulary, which is high compared to a smaller $\delta$ value such as $2.0$.
For the Exponential Minimum Sampling or EXP as shown in Figure  8(b), all languages are closely stable in terms of the curve characteristics.
Method Language $\gamma$ Metrics ($z=4.0$) Metrics ($z=5.0$)
TPR TNR FPR FNR TPR TNR FPR FNR
KGW English 0.1 0.998 1.0 0.0 0.002 0.994 1.0 0.0 0.006
0.5 0.998 1.0 0.0 0.002 0.998 1.0 0.0 0.002
0.9 0.162 1.0 0.0 0.838 0.0 1.0 0.0 1.0
Arabic 0.1 0.970 0.990 0.010 0.030 0.954 0.994 0.006 0.046
0.5 0.974 0.994 0.006 0.026 0.958 0.998 0.002 0.042
0.9 0.112 1.0 0.0 0.888 0.0 1.0 0.0 1.0
Chinese 0.1 1.0 0.992 0.008 0.0 0.994 0.992 0.008 0.006
0.5 1.0 0.998 0.002 0.0 0.992 1.0 0.0 0.008
0.9 0.162 1.0 0.0 0.838 0.0 1.0 0.0 1.0
Indonesian 0.1 0.956 1.0 0.0 0.044 0.902 1.0 0.0 0.098
0.5 0.966 0.996 0.004 0.034 0.862 1.0 0.0 0.138
0.9 0.026 1.0 0.0 0.974 0.0 1.0 0.0 1.0
Unigram English 0.1 0.986 0.998 0.002 0.014 0.936 1.0 0.0 0.064
0.5 1.0 0.990 0.010 0.0 1.0 0.998 0.002 0.0
0.9 0.244 1.0 0.0 0.756 0.0 1.0 0.0 1.0
Arabic 0.1 0.976 0.970 0.030 0.024 0.948 0.976 0.024 0.052
0.5 0.986 0.970 0.030 0.014 0.970 0.996 0.004 0.030
0.9 0.432 0.998 0.002 0.568 0.0 1.0 0.0 1.0
Chinese 0.1 1.0 0.856 0.144 0.0 1.0 0.926 0.074 0.0
0.5 0.998 0.876 0.124 0.002 0.996 0.926 0.074 0.004
0.9 0.500 0.946 0.054 0.500 0.0 1.0 0.0 1.0
Indonesian 0.1 0.952 0.970 0.030 0.048 0.922 0.982 0.018 0.078
0.5 0.984 0.954 0.046 0.016 0.962 0.990 0.010 0.038
0.9 0.052 0.998 0.002 0.948 0.0 1.0 0.0 1.0
5 : Performance Metrics for Multiple Languages and Watermarking Methods calculated by following the test statistics score with two threshold values as shown by $z$. The results show the effect of varying $\gamma$ values with a fixed $\delta=2.0$ for KGW Kirchenbauer et al. (2023a) and Unigram Zhao et al. (2023).
Method Language $\delta$ Metrics ($z=4.0$) Metrics ($z=5.0$)
TPR TNR FPR FNR TPR TNR FPR FNR
KGW English 2.0 0.998 1.0 0.0 0.002 0.998 1.0 0.0 0.002
5.0 1.0 1.0 0.0 0.0 1.0 1.0 0.0 0.0
10.0 1.0 1.0 0.0 1.0 1.0 1.0 0.0 0.0
Arabic 2.0 0.974 0.994 0.006 0.026 0.958 0.998 0.002 0.042
5.0 1.0 0.998 0.002 0.0 1.0 1.0 0.0 0.0
10.0 1.0 0.998 0.002 0.0 1.0 1.0 0.0 0.0
Chinese 2.0 1.0 0.998 0.002 0.0 0.992 1.0 0.0 0.008
5.0 1.0 0.994 0.006 0.0 1.0 0.998 0.002 0.0
10.0 1.0 0.998 0.002 0.0 1.0 1.0 0.0 0.0
Indonesian 2.0 0.966 0.996 0.004 0.034 0.862 1.0 0.0 0.138
5.0 0.996 1.0 0.0 0.004 0.992 1.0 0.0 0.008
10.0 1.0 1.0 0.0 0.0 1.0 1.0 0.0 0.0
Unigam English 2.0 1.0 0.990 0.010 0.0 1.0 0.998 0.002 0.0
5.0 1.0 0.984 0.016 0.0 1.0 1.0 0.0 0.0
10.0 1.0 0.984 0.016 0.0 1.0 1.0 0.0 0.0
Arabic 2.0 0.986 0.970 0.030 0.014 0.97 0.996 0.004 0.03
5.0 1.0 0.972 0.028 0.0 1.00 0.990 0.010 0.00
10.0 1.0 0.960 0.040 0.0 1.00 0.990 0.010 0.00
Chinese 2.0 0.998 0.876 0.124 0.002 0.996 0.926 0.074 0.004
5.0 1.0 0.890 0.110 0.0 1.0 0.924 0.076 0.0
10.0 1.0 0.878 0.122 0.0 1.0 0.922 0.078 0.0
Indonesian 2.0 0.984 0.954 0.046 0.016 0.962 0.990 0.010 0.038
5.0 1.0 0.962 0.038 0.0 1.0 0.994 0.006 0.0
10.0 1.0 0.942 0.058 0.0 1.0 0.984 0.016 0.0
Metric ($z=0.2$) Metric ($z=0.3$)
TPR TNR FPR FNR TPR TNR FPR FNR
XSIR English 2.0 1.0 0.702 0.298 0.0 0.994 0.958 0.042 0.006
5.0 1.0 0.734 0.266 0.0 0.998 0.964 0.036 0.002
10.0 0.996 0.708 0.292 0.004 0.996 0.960 0.040 0.004
Arabic 2.0 0.998 0.916 0.084 0.002 0.996 0.994 0.006 0.004
5.0 1.0 0.928 0.072 0.0 0.998 0.994 0.006 0.002
10.0 1.0 0.898 0.102 0.0 0.998 0.982 0.018 0.002
Chinese 2.0 1.0 0.966 0.034 0.0 0.994 0.994 0.006 0.006
5.0 0.996 0.944 0.056 0.004 0.992 0.982 0.018 0.008
10.0 0.992 0.964 0.036 0.008 0.992 0.994 0.006 0.008
Indonesian 2.0 0.988 0.990 0.010 0.012 0.976 0.998 0.002 0.024
5.0 0.998 0.994 0.006 0.002 0.998 1.0 0.0 0.002
10.0 0.996 0.984 0.016 0.004 0.994 0.998 0.002 0.006
Metric ($p=10e-5$) Metric ($p=10e-4$)
TPR TNR FPR FNR TPR TNR FPR FNR
EXP English - 0.998 1.0 0.0 0.002 0.998 0.996 0.004 0.002
Arabic - 0.982 1.0 0.0 0.018 0.982 1.0 0.0 0.018
Chinese - 0.994 1.0 0.0 0.006 0.994 1.0 0.0 0.006
Indonesian - 0.990 1.0 0.0 0.010 0.992 1.0 0.0 0.008
6 : Performance Metrics for Multiple Languages and Watermarking Methods calculated by following the test statistics score. The results show the effect of varying $\delta$ values with a fixed of $\gamma=0.5$ for KGW, Unigram, and XSIR. Using a threshold of $0.2$ for XSIR resulted in higher FPRs Specifically for the English language. (see note below).
(5)(5)footnotetext:
XSIR logically employs 50-50 split ratio. It’s not a hyper-parameter that can be changed directly.
(a) Translation-¿Paraphrase Attacks
(b) Translation-¿Paraphrase-¿Translation Attacks
9 : Watermark detection ROC curves with AUC after attacks for syntactical methods KGW, Unigram, and EXP. We fix $\gamma=0.5$ and $\delta=2.0$ for KGW and Unigram. For each attack, the watermark threshold is calculated automatically by comparing unwatermarked and attacked watermarked text score for $100$ generations. The bottom row in all figures represent XSIR original attack setting in which non-English language is used first.
2.2 Watermark Detection After Attacks for Syntactical Methods
In Figure 9(b) shows the completion of attack pipeline performed on KGW, Unigram, and EXP. Translation followed by paraphrase attacks are no worse than translations alone for syntactical methods. However, we notice that beginning with English in the pipeline performs better in all methods as can be seen from the top row of all the figures. Translation-paraphrase-translation performs well when English is the source language in all methods across all target languages, whereas the detection is very low when English is not the source language.